# Category $\mathcal{O}$ for hybrid quantum groups and non-commutative Springer resolutions

### v3

@@ 7.2.1. Grading on a category
7.2.2. Graded ${\mathrm{A}}$-modules
Let $P(\lambda)_{\mathbb{C}}$ be the projective cover of $L(\lambda)_{\mathbb{C}}$ in $u_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$. Let $Z(\lambda)_{\mathbb{C}}=u_{\zeta}\otimes_{u_{\zeta}^{\geq 0}}{\mathbb{C}}_{\lambda}$ be the baby Verma module. Recall that $P^{b}(\lambda)_{\mathbb{C}}$ is the project cover of $L(\lambda)_{\mathbb{C}}$ in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$. We have isomorphisms of ${\mathfrak{U}}^{b}_{\zeta}$-modules
@@ 7.2.2. Graded ${\mathrm{A}}$-modules
where the left horizontal maps are by forgetful functors, and the right inclusions are induced by the quotients ${\mathrm{A}}_{\mathfrak{n}}\twoheadrightarrow{\mathrm{A}}_{0}$ and ${\mathfrak{U}}^{b}_{\zeta}\twoheadrightarrow u_{\zeta}$. For $x\in W_{\mathrm{ex}}$, we denote by ${\mathcal{L}}_{x}$, ${\mathcal{Z}}_{x}$, ${\mathcal{P}}_{x}$, ${\mathcal{P}}^{b}_{x}$, ${{\mathcal{Q}}}_{x}^{\leq\nu}$ the preimages of $L(x\bullet_{l}0)_{\mathbb{C}}$, $Z(x\bullet_{l}0)_{\mathbb{C}}$, $P(x\bullet_{l}0)_{\mathbb{C}}$, $P^{b}(x\bullet_{l}0)_{\mathbb{C}}$, $Q(x\bullet_{l}0)^{\leq l\nu}_{\mathbb{C}}$ under the vertical equivalences above. By Lemma 2.2, the ${\mathrm{A}}_{\mathfrak{n}}$-module ${\mathcal{L}}_{x}$ can be endowed with a $B$-equivariant structure, which makes ${\mathcal{L}}_{x}$ a simple object in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$. Recall the Verma objects ${\mathcal{M}}_{x}$ ($x\in W_{\mathrm{ex}}$) in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$. Set ${\mathcal{M}}^{\mathbb{C}}_{x}={\mathcal{M}}_{x}\otimes_{S}{\mathbb{C}}$. We call ${\mathcal{M}}^{\mathbb{C}}_{x}$ (resp. ${\mathcal{Z}}_{x}$) the Verma object in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$ and ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}$ (resp. in ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T}$). A Verma flag is a filtration with composition factors given by Verma objects.
Lemma 7.3.
Let $x\in W_{\mathrm{ex}}$.
  1. (1)
    The objects ${\mathcal{Z}}_{x}$, ${\mathcal{P}}_{x}$, ${\mathcal{P}}^{b}_{x}$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}$ admit liftings $\widetilde{{\mathcal{Z}}}_{x}$, $\widetilde{{\mathcal{P}}}_{x}$, $\widetilde{{\mathcal{P}}}^{b}_{x}$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$, and the liftings are unique up to grading shifts.
  2. (2)
    The objects ${\mathcal{L}}_{x}$, ${\mathcal{M}}^{\mathbb{C}}_{x}$, ${\mathcal{Q}}_{x}^{\leq\nu}$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$ admits liftings ~$\widetilde{{\mathcal{L}}}_{x}$, ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}$, $\widetilde{{\mathcal{Q}}}_{x}^{\leq\nu}$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$, and the liftings are unique up to grading shifts.
  3. (3)
    Let $P$ be an object in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$ with a lifting $\widetilde{P}$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$. If $P$ admits a Verma flag, then $\widetilde{P}$ also admits a Verma flag (i.e. a filtration with composition factors of the form ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle$). Similar result holds for ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}$ and ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T}$.
Proof.
By a standard argument (see e.g. [12]), one can show that any $B$-equivariant simple ${\mathrm{A}}_{\mathfrak{n}}$-module (resp. $T$-equivariant projective module of ${\mathrm{A}}_{\mathfrak{n}}$ or ${\mathrm{A}}_{0}$) can be equipped with a ${\mathbb{C}}^{\times}$-equivariant structure, and up to isomorphism any lifting differs by a ${\mathbb{C}}^{\times}$-character.
By its construction (4.15), any Verma object ${\mathcal{M}}^{\mathbb{C}}_{x}$ can be naturally equipped with a ${\mathbb{C}}^{\times}$-equivariant structure. Let ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}$ be a lifting of ${\mathcal{M}}^{\mathbb{C}}_{x}$, then $\widetilde{{\mathcal{Z}}}_{x}=\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}{\mathbb{C}}$ provides a lifting of ${\mathcal{Z}}_{x}$ by (7.4). Suppose ~$\widetilde{{\mathcal{M}}}^{{\mathbb{C}},1}_{x}$, ~$\widetilde{{\mathcal{M}}}^{{\mathbb{C}},2}_{x}$ are two liftings of ${\mathcal{M}}^{\mathbb{C}}_{x}$. Since $\mathrm{End}({\mathcal{M}}^{\mathbb{C}}_{x})={\mathbb{C}}$, by (7.3) we have $\mathrm{Hom}(\widetilde{{\mathcal{M}}}^{{\mathbb{C}},1}_{x},\widetilde{{\mathcal{M}}}^{{\mathbb{C}},2}_{x}\langle i\rangle)={\mathbb{C}}\delta_{i,d}$ for some $d\in{\mathbb{Z}}$, and any nonzero map from ~$\widetilde{{\mathcal{M}}}^{{\mathbb{C}},1}_{x}$ to ~$\widetilde{{\mathcal{M}}}^{{\mathbb{C}},2}_{x}\langle d\rangle$ is an isomorphism. The uniqueness of $\widetilde{{\mathcal{Z}}}_{x}$ follows similarly. This proves (1) as well as the assertion for ${\mathcal{L}}_{x}$ and ${\mathcal{M}}^{\mathbb{C}}_{x}$ in (2).
@@
Now we show (3). We only show the assertion for ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$. We prove by increasing induction on the length of Verma flag $\sum_{x}(P:{\mathcal{M}}^{\mathbb{C}}_{x})$. Let $x$ be maximal (with respect to the order $\leq^{\frac{\infty}{2}}$) among the elements $y\in W_{\mathrm{ex}}$ such that ${\mathcal{M}}^{\mathbb{C}}_{y}$ is a Verma factor of $P$. Then we have $\mathrm{Hom}(\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle,\widetilde{P})\neq 0$ for some $d$, and any nonzero element gives an inclusion ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle\hookrightarrow\widetilde{P}$ (since any nonzero map ${\mathcal{M}}^{\mathbb{C}}_{x}\rightarrow P$ is an inclusion). Now $\widetilde{P}/\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle$ is a lifting of $P/{\mathcal{M}}^{\mathbb{C}}_{x}$, which by induction hypothesis admits a Verma flag. Hence $\widetilde{P}$ admits a Verma flag.
Finally, we show the assertion for ${\mathcal{Q}}_{x}^{\leq\nu}$ in (2). Consider the ${\mathbb{C}}^{\times}$-equivariant $({\mathrm{A}}_{n}\rtimes U{\mathfrak{n}})$-module $\widetilde{{\mathcal{Q}}}_{x}=({\mathrm{A}}_{n}\rtimes U{\mathfrak{n}})\otimes_{{\mathrm{A}}_{\mathfrak{n}}}\widetilde{{\mathcal{P}}}^{b}_{x}$. By (3), $\widetilde{{\mathcal{P}}}^{b}_{x}$ admits a Verma flag. Hence by Lemma 5.6, $\widetilde{{\mathcal{Q}}}_{x}$ also admits a Verma flag. By (the non-defomed version of) Proposition 5.5, we have $\mathrm{Ext}^{1}(\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y},\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y^{\prime}}\langle d\rangle)\neq 0$ only if $y\leq^{\frac{\infty}{2}}y^{\prime}$. We define $\widetilde{{\mathcal{Q}}}_{x}^{\leq\nu}$ as the quotient of $\widetilde{{\mathcal{Q}}}_{x}$ by the submodule composed by ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{t(\mu)y}$ with $\mu\nleq\nu$ and $y\in W$. As in Lemma 2.3(2), $\widetilde{{\mathcal{Q}}}_{x}^{\leq\nu}$ is a lifting of ${\mathcal{Q}}_{x}^{\leq\nu}$. For the uniqueness of the lifting, we set ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times},\leq\nu}$ as the Serre subcategory of ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$ generated by ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{t(\mu)y}$ with $\mu\leq\nu$ and $y\in W$. Then any lifting of ${\mathcal{Q}}_{x}^{\leq\nu}$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$ is a projective cover of ~$\widetilde{{\mathcal{L}}}_{x}\langle d\rangle$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times},\leq\nu}$ for some $d$, which thus is unique up to a grading shift. ∎
@@
For an object $P$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$ admitting Verma flags, we set $(P:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle)$ as the multiplicity of the Verma factor ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle$, which does not depend on the choice of Verma flag. We define the graded multiplicity as
@@
For any object $M$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$, it admits a separated decreasing ${\mathbb{Z}}_{\leq 0}$-filtration (usually of infinite length) with simple composition factors. We set $[M:\widetilde{{\mathcal{L}}}_{x}\langle d\rangle]$ as the multiplicity (which is finite) of ~$\widetilde{{\mathcal{L}}}_{x}\langle d\rangle$ in the factors, which dose not depend on the choice of such filtration, and we define the graded multiplicity
(7.6) $[M:\widetilde{{\mathcal{L}}}_{x}]_{{\mathrm{gr}}}=\sum_{d}[M:\widetilde{{\mathcal{L}}}_{x}\langle d\rangle]\cdot v^{d}\ \in{\mathbb{Z}}[v^{\pm 1}].$
We also introduce similar notations for ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$ and ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$.
Let $\hat{K}_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$ be the space of ${\mathbb{Z}}[v^{\pm 1}]$-linear formal series in {[~$\{[\widetilde{{\mathcal{L}}}_{x}]\}_{x}$ that are bounded above (with respect to the order $\leq^{\frac{\infty}{2}}$). We have an inclusion
$K_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})\rightarrow\hat{K}_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}),\quad[M]\mapsto\sum_{x}[M:\widetilde{{\mathcal{L}}}_{x}]_{{\mathrm{gr}}}\cdot[\widetilde{{\mathcal{L}}}_{x}].$
Note that (the image of) {[~$\{[\widetilde{{\mathcal{M}}}_{x}]\}_{x}$, $\{[\widetilde{{\mathcal{Z}}}_{x}]\}_{x}$ give other topological basis on $\hat{K}_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$.

### v4

@@ 7.2.1. Grading on a category
7.2.2. Graded $\mathsf{A}$}-modules
Let $P(\lambda)_{\mathbb{C}}$ be the projective cover of $L(\lambda)_{\mathbb{C}}$ in $u_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$. Let $Z(\lambda)_{\mathbb{C}}=u_{\zeta}\otimes_{u_{\zeta}^{\geq 0}}{\mathbb{C}}_{\lambda}$ be the baby Verma module. Recall that $P^{b}(\lambda)_{\mathbb{C}}$ is the projective cover of $L(\lambda)_{\mathbb{C}}$ in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$. We have isomorphisms of ${\mathfrak{U}}^{b}_{\zeta}$-modules
@@ 7.2.2. Graded $\mathsf{A}$}-modules
(7.5) ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}$${\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T}$${\mathcal{O}}^{0}_{\mathbb{C}}$${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{\mathbb{C}}$$u_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{\mathbb{C}},$$\scriptstyle{\simeq}\scriptstyle{\simeq}\scriptstyle{\simeq}$
where the left horizontal maps are by forgetful functors, and the right inclusions are induced by the quotients ${\mathsf{A}}_{\mathfrak{n}}\twoheadrightarrow{\mathsf{A}}_{0}$ and ${\mathfrak{U}}^{b}_{\zeta}\twoheadrightarrow u_{\zeta}$. For any $x\in W_{\mathrm{ex}}$ and bounded above poset ideal $\Omega\subset\Lambda$, we denote by
${\mathcal{L}}_{x},\quad{\mathcal{Z}}_{x},\quad{\mathcal{P}}_{x},\quad{\mathcal{P}}^{b}_{x},\quad{{\mathcal{Q}}}_{x}^{\Omega}$
the preimages of $L(x\bullet_{l}0)_{\mathbb{C}}$, $Z(x\bullet_{l}0)_{\mathbb{C}}$, $P(x\bullet_{l}0)_{\mathbb{C}}$, $P^{b}(x\bullet_{l}0)_{\mathbb{C}}$, $Q(x\bullet_{l}0)^{\Omega}_{\mathbb{C}}$ under the equivalences above. By Lemma 2.3, the ${\mathsf{A}}_{\mathfrak{n}}$-module ${\mathcal{L}}_{x}$ can be endowed with a $B$-equivariant structure, which makes ${\mathcal{L}}_{x}$ a simple object in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$. Recall the Verma objects ${\mathcal{M}}_{x}$ ($x\in W_{\mathrm{ex}}$) in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$. We set
${\mathcal{M}}^{\mathbb{C}}_{x}={\mathcal{M}}_{x}\otimes_{S}{\mathbb{C}}$
and call ${\mathcal{M}}^{\mathbb{C}}_{x}$ (resp. ${\mathcal{Z}}_{x}$) the Verma object in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$ and ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}$ (resp. in ${\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T}$). A Verma flag is a finite filtration with composition factors given by Verma objects.
Lemma 7.3.
Let $x\in W_{\mathrm{ex}}$, and let $\Omega\subset\Lambda$ be a bounded above poset ideal.
  1. (1)
    The objects ${\mathcal{Z}}_{x}$, ${\mathcal{P}}_{x}$, ${\mathcal{P}}^{b}_{x}$ in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}$ admit liftings $\widetilde{{\mathcal{Z}}}_{x}$, $\widetilde{{\mathcal{P}}}_{x}$, $\widetilde{{\mathcal{P}}}^{b}_{x}$ in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$, and the liftings are unique up to grading shifts.
  2. (2)
    The objects ${\mathcal{L}}_{x}$, ${\mathcal{M}}^{\mathbb{C}}_{x}$, ${\mathcal{Q}}_{x}^{\Omega}$ in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$ admits liftings $\widetilde{{\mathcal{L}}}_{x}$, $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}$, $\widetilde{{\mathcal{Q}}}_{x}^{\Omega}$ in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$, and the liftings are unique up to grading shifts.
  3. (3)
    Let $P$ be an object in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$ with a lifting $\widetilde{P}$ in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$. If $P$ admits a Verma flag, then $\widetilde{P}$ also admits a Verma flag (i.e. a filtration with composition factors of the form $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle$). Similar result holds for ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}$ and ${\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T}$.
Proof.
@@
[…]
@@
Step 2. Now we show (3). We only show the assertion for ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$. We prove by increasing induction on the length of Verma flag $\sum_{x}(P:{\mathcal{M}}^{\mathbb{C}}_{x})$. Let $x$ be such that $x\bullet_{l}0$ is maximal among the weights $y\bullet_{l}0$ with ${\mathcal{M}}^{\mathbb{C}}_{y}$ appearing as a Verma factor of $P$. Then we have $\mathrm{Hom}(\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle,\widetilde{P})\neq 0$ for some $d$, and any nonzero element gives an inclusion $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle\hookrightarrow\widetilde{P}$ (since any nonzero map ${\mathcal{M}}^{\mathbb{C}}_{x}\rightarrow P$ is an inclusion). Now $\widetilde{P}/\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle$ is a lifting of $P/{\mathcal{M}}^{\mathbb{C}}_{x}$, which by induction hypothesis admits a Verma flag. Hence $\widetilde{P}$ admits a Verma flag.
Step 3. We show the assertion for ${\mathcal{Q}}_{x}^{\Omega}$ in (2). Recall the truncated categories ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T,\Omega}$, ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{B,\Omega}$, and the truncation functors $\tau^{T,\Omega}$, $\tau^{U{\mathfrak{n}},\Omega}$ introduced in §5.2.2. We set ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T}={\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T}\cap{\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T,\Omega}$ and define ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B,\Omega}$ similarly. The truncation functors restrict to
$\tau^{T,\Omega}:\ {\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T}\rightarrow{\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T,\Omega}\quad\text{and}\quad\tau^{U{\mathfrak{n}},\Omega}:\ ({\mathsf{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}\rightarrow{\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B,\Omega}.$
We set ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T\times{\mathbb{C}}^{\times},\Omega}$ as the preimage of ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T,\Omega}$ under the degrading functor ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T\times{\mathbb{C}}^{\times}}\rightarrow{\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T}$, and define ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B\times{\mathbb{C}}^{\times},\Omega}$ similarly. Consider the truncation functor
$\tau^{T\times{\mathbb{C}}^{\times},\Omega}:\ {\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T\times{\mathbb{C}}^{\times}}\rightarrow{\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T\times{\mathbb{C}}^{\times},\Omega},$
by sending $M$ to its quotient by the minimal graded submodule containing $\ker(M\rightarrow\tau^{T,\Omega}(M))$. The functor $\tau^{T\times{\mathbb{C}}^{\times},\Omega}$ is left adjoint to the natural inclusion. We define
$\tau^{U{\mathfrak{n}}\times{\mathbb{C}}^{\times},\Omega}:\ ({\mathsf{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T\times{\mathbb{C}}^{\times}}\rightarrow{\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B\times{\mathbb{C}}^{\times},\Omega},$
by sending $M$ to its quotient by the (graded) ${\mathsf{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}}$-module generated by $\ker(M\rightarrow\tau^{T\times{\mathbb{C}}^{\times},\Omega}(M))$. Then $\tau^{U{\mathfrak{n}}\times{\mathbb{C}}^{\times},\Omega}$ is left adjoint to the natural inclusion.
We claim that the following natural transformations are isomorphisms
(7.6) $\tau^{T,\Omega}\circ\upsilon{\ \xrightarrow{\sim}\ }\upsilon\circ\tau^{T\times{\mathbb{C}}^{\times},\Omega}\quad\text{and}\quad\tau^{U{\mathfrak{n}},\Omega}\circ\upsilon{\ \xrightarrow{\sim}\ }\upsilon\circ\tau^{U{\mathfrak{n}}\times{\mathbb{C}}^{\times},\Omega}$
where $\upsilon$ is the degrading functor. Note that the second isomorphism follows from the first one, hence we only need to show the first isomorphism. By the right exactness of truncation functors and degrading functors, it is enough to construct isomorphism $\upsilon\circ\tau^{T\times{\mathbb{C}}^{\times},\Omega}(P^{\mathrm{gr}})\simeq\tau^{T,\Omega}(P)$ for any $T$-equivariant projective ${\mathsf{A}}_{\mathfrak{n}}$-module $P$ and its lifting $P^{\mathrm{gr}}$ (which exists by (1)). By (3), $P^{\mathrm{gr}}$ admits a Verma filtration. By (2.4) and (7.3) applied to the degrading $D^{\mathrm{b}}({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}})\rightarrow D^{\mathrm{b}}({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T})$, we have $\mathrm{Ext}^{1}(\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x},\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\langle d\rangle)=0$ if $x\bullet_{l}0\nless y\bullet_{l}0$. Hence $\tau^{T\times{\mathbb{C}}^{\times},\Omega}(P^{\mathrm{gr}})$ is the quotient of $P^{\mathrm{gr}}$ by the submodule composed by Verma factors $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\langle d\rangle$ with $y\bullet_{l}0\notin\Omega$, which shows that $\tau^{T\times{\mathbb{C}}^{\times},\Omega}(P^{\mathrm{gr}})=\tau^{T,\Omega}(P)$ after forgetting the grading structure. It proves the claim.
Finally, consider the induction functor
$\mathrm{ind}_{c}:=({\mathsf{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}})\otimes_{{\mathsf{A}}_{\mathfrak{n}}}-:\ {\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}\rightarrow({\mathsf{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}.$
Let $x\in W_{\mathrm{ex}}$ such that $x\bullet_{l}0\in\Omega$. The module $\widetilde{{\mathcal{Q}}}_{x}^{\Omega}=\tau^{U{\mathfrak{n}}\times{\mathbb{C}}^{\times},\Omega}(\mathrm{ind}_{c}\widetilde{{\mathcal{P}}}^{b}_{x})$ is the projective cover of $\widetilde{{\mathcal{L}}}_{x}\langle d\rangle$ (for some $d\in{\mathbb{Z}}$) in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times},\Omega}$, which provide a lifting of ${\mathcal{Q}}^{\Omega}_{x}=\tau^{U{\mathfrak{n}},\Omega}(\mathrm{ind}_{c}{{\mathcal{P}}}^{b}_{x})$. The uniqueness of lifting follows from the uniqueness of projective cover. ∎
@@
For an object $P$ in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$ admitting Verma flags, we set $(P:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle)$ as the multiplicity of the Verma factor $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle$, which does not depend on the choice of Verma flag. We define the graded multiplicity as
@@
For any object $M$ in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$, it admits a separated decreasing ${\mathbb{Z}}_{\leq 0}$-filtration (may of infinite length) with simple composition factors. We set $M:\widetilde{{\mathcal{L}}}_{x}\langle d\rangle]$ as the multiplicity (which is finite) of $\widetilde{{\mathcal{L}}}_{x}\langle d\rangle$ in the factors, which dose not depend on the choice of such filtration, and we define the graded multiplicity
(7.8) $M:\widetilde{{\mathcal{L}}}_{x}]_{{\mathrm{gr}}}=\sum_{d}[M:\widetilde{{\mathcal{L}}}_{x}\langle d\rangle]\cdot v^{d}\ \in{\mathbb{Z}}[v^{\pm 1}].$
We also introduce similar notations for ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$ and ${\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$.
Let $\hat{K}_{0}({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$ be the space of ${\mathbb{Z}}[v^{\pm 1}]$-linear formal series in $\{[\widetilde{{\mathcal{L}}}_{x}]\}_{x}$ that are bounded above (with respect to the order $\leq^{\frac{\infty}{2}}$). We have an inclusion
$K_{0}({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})\rightarrow\hat{K}_{0}({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}),\quad[M]\mapsto\sum_{x}[M:\widetilde{{\mathcal{L}}}_{x}]_{{\mathrm{gr}}}\cdot[\widetilde{{\mathcal{L}}}_{x}].$
Note that (the images of) $\{[\widetilde{{\mathcal{M}}}_{x}]\}_{x}$ and $\{[\widetilde{{\mathcal{Z}}}_{x}]\}_{x}$ give other topological basis on $\hat{K}_{0}({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$.